Let $x$ and $y$ be integers. If $x$ and $y$ satisfy $41x + 5y = 31$, then find the residue of $x$ modulo $5$.
Answer by ikleyn(52775)

.
Let x and y be integers. If x and y satisfy 41x + 5y = 31, then find the residue of x modulo 5.
~~~~~~~~~~~~~~~~~~~

You are given this equation

    41x + 5y = 31,    (1)

where x and y are integers.  Hence

    41x = 31 - 5y.    (2)


They want you find residue  {x mod 5} from modular equation (2).


It means that you consider equation (2)  as an equivalence by modulo 5.
If so, you may forget about the term -5y, because it is 0 (zero) by the modulo 5.


So, your task is to find the residue {x mod 5} from the modular equation

    41x = 31 mod 5,

which is equivalent to 

    41x = 1  mod 5.    (3)


The set of possible solutions is  x = {0 mod 5}, {1 mod 5}, {2 mod 5}, {3 mod 5} and {4 mod 5}.


    x = {0 mod 5} turns (3)  into  41*0 = 1 mod 5,  which is obviously wrong. 

    x = {1 mod 5} turns (3)  into  41*1 = 1 mod 5,  which is obviously correct modular equation ( since 41 mod 5 = 1 mod 5). 

    x = {2 mod 5} turns (3)  into  41*2 = 1 mod 5,  which is obviously wrong. 

    x = {3 mod 5} turns (3)  into  41*3 = 1 mod 5,  which is obviously wrong. 

    x = {4 mod 5} turns (3)  into  41*4 = 1 mod 5,  which is obviously wrong. 


Thus of five possible solutions only one is valid:  x = {1 mod 5}.    ANSWER


So, in this problem, you can easily find the answer by the brute force method.

The problem can be solved using more refined methods, but this reasoning is the number 1,
which you should know and understand as a basic method to start thinking.